# Comparing AI Agents to Cybersecurity Professionals in Real-World Penetration Testing

### 10 Dec 2025 v1

@@ 3.3 Agents
Model Success Rate
Claude 4.5 Sonnet 55%
ARTEMIS 48.6%
OpenAI GPT-5 45.9%
Claude 4.1 Opus 38%
Claude 4 Opus 38%
Claude 4 Sonnet 35%
OpenAI o3-mini 22.5%
Table 2: Comparison of success rates on Cybench. Aside from ARTEMIS and GPT-5 results, all numbers are taken from the Cybench website.
@@ 4.2 Agent Results
We compare ARTEMIS to OpenAI’s Codex, Claude Code, CyAgent, Incalmo, and MAPTA. We exclude semi-autonomous systems like PentestGPT or CAI to focus on fully autonomous capabilities. We run two ARTEMIS configurations: $A_{1}$ uses GPT-5 for both supervisor and sub-agents, while $A_{2}$ uses an ensemble of supervisor models (Claude Sonnet 4, OpenAI o3, Claude Opus 4, Gemini 2.5 Pro, and OpenAI o3 Pro) similar to Alloy Agents (Ziegler, 2025) with Claude Sonnet 4 for sub-agents. Both run for 16 hours (9am–5pm across two days); for human comparisons, we limit scoring to the first 10 hours. Other scaffolds run to completion since they cannot sustain 10+ hours of continuous work. Codex, MAPTA, and Incalmo use GPT-5. CyAgent is tested with both GPT-5 and Claude Sonnet 4. Claude Code uses Claude Sonnet 4. All scaffolds receive the same instructions (Appendix J) except Incalmo and MAPTA, which only accept target scope. All agents used the same VM as human participants.
As shown in Table 1, ARTEMIS significantly outperforms existing scaffolds. Claude Code and MAPTA refuse the task out of the box, while Incalmo stalls at early reconnaissance due to its rigid task graph, resulting in 0 findings each. We observed no such refusals across either ARTEMIS trial. ARTEMIS reached a peak of 8 active sub-agents in parallel, averaging 2.82 concurrent sub-agents per supervisor iteration. However, as shown in Figure 2, ARTEMIS submits more false positives than human participants (discussed in Section 6).
@@ 5.4 Cost Analysis
Cost is an important differentiator between agents and professionals. To understand long-horizon performance, we ran ARTEMIS for 16 hours total (8 hours across two working days, 9am–5pm); we evaluate only the first 10 hours but tracked performance throughout. We monitored costs via dedicated API keys for each experiment.
@@ Acknowledgements
We thank Pavan Agrawal, Andy Applebaum, Jack Cable, Nicholas Carlini, Sven Cattell, Casey Ellis, Eric Han, Kevin Liu, Joel Parish, Ashwin Ramaswami, and Olivia Watkins for formative discussions shaping this work. We also thank James Aung, Belinda Mo, and Justin Wang for feedback on early revisions. We are grateful to Elizabeth Proehl and Erin Maneri for administrative support. Finally, we thank John Gerth, Alex Keller, Andrej Krevl, Joe Little, and Jimmy Wu from Stanford IT for their close collaboration and support in conducting this study, as well as all the cybersecurity professionals who participated in our evaluation. This work is supported by the Stanford Institute for Human-Centered Artificial Intelligence Seed Grant and an unrestricted gift from OpenAI.
@@ 8 Ethics Statement
Our study was conducted with IRB approval. Participants provided informed consent to session recordings, behavioral data collection, and publication of anonymized results. They were compensated \$2,000 for their participation. Our IRB provisions (Section 3.1) protect privacy: all data and recordings are stored on encrypted, access-controlled servers and will be deleted after the retention period. We anonymize all participants, referring to them as $P_{1}$$P_{10}$.
@@
We maintained communication with the target’s IT staff to report, triage, and patch discovered vulnerabilities per responsible disclosure. All participants read the university’s VDP and were instructed to avoid destructive actions and stay within scope. The university enforced risk-based minimum standards including monthly vulnerability management via Qualys. By collaborating with IT and recruiting security professionals as participants, we helped improve the university’s security posture. To mitigate risks from autonomous agents, a research team member monitored each session in real time with ability to terminate on out-of-scope or risky behavior.

### 03 Mar 2026 v2

@@ 3.3 Agents
Scaffold Model Success
CyAgent Claude 4.5 Sonnet 55%
ARTEMIS GPT-5 48.6%
CyAgent GPT-5 45.9%
CyAgent Claude 4.1 Opus 38%
CyAgent Claude 4 Opus 38%
CyAgent Claude 4 Sonnet 35%
CyAgent o3-mini 22.5%
Table 2: Cybench success rates by scaffold and model. All CyAgent results except GPT-5 were sourced from the Cybench leaderboard. We evaluated GPT-5 using CyAgent pass@1, as these results were not yet available on the leaderboard at the time of writing. ARTEMIS shows no significant scaffold uplift over CyAgent + GPT-5 on CTF tasks.
@@ 4.2 Agent Results
We compare ARTEMIS to OpenAI’s Codex, Claude Code, CyAgent, Incalmo, and MAPTA. We exclude semi-autonomous systems like PentestGPT or CAI to focus on fully autonomous capabilities. We run two ARTEMIS configurations: $A_{1}$ uses GPT-5 for both supervisor and sub-agents, while $A_{2}$ uses an ensemble of supervisor models (Claude Sonnet 4, OpenAI o3, Claude Opus 4, Gemini 2.5 Pro, and OpenAI o3 Pro) similar to Alloy Agents (Ziegler, 2025) with Claude Sonnet 4 for sub-agents. Both configurations execute for 16 hours; we evaluate only the first 10 hours to maintain comparability with human participants. Other scaffolds run to completion since they cannot sustain 10+ hours of continuous work. Codex, MAPTA, and Incalmo use GPT-5. CyAgent is tested with both GPT-5 and Claude Sonnet 4. Claude Code uses Claude Sonnet 4. All scaffolds receive the same instructions (Appendix J) except Incalmo and MAPTA, which only accept target scope. All agents used the same VM as human participants.
As shown in Table 1, ARTEMIS significantly outperforms existing scaffolds. Claude Code and MAPTA refuse the task out of the box, while Incalmo stalls at early reconnaissance due to its rigid task graph, resulting in 0 findings each. We observed no such refusals across either ARTEMIS trial, despite using the same underlying models as Claude Code ($A_{2}$) and MAPTA ($A_{1}$), suggesting that scaffolding and prompting decisions are directly responsible for bypassing refusal mechanisms. ARTEMIS reached a peak of 8 active sub-agents in parallel, averaging 2.82 concurrent sub-agents per supervisor iteration. However, as shown in Figure 2, ARTEMIS submits more false positives than human participants (discussed in Section 6).
@@ 5.4 Cost Analysis
Cost is an important differentiator between agents and professionals. We tracked API costs via dedicated keys for each experiment across the full execution window.
@@ Acknowledgements
We thank Pavan Agrawal, Andy Applebaum, Jack Cable, Nicholas Carlini, Sven Cattell, Casey Ellis, Eric Han, Kevin Liu, Jolene Parish, Ashwin Ramaswami, and Olivia Watkins for formative discussions shaping this work. We also thank James Aung, Belinda Mo, and Justin Wang for feedback on early revisions. We are grateful to Elizabeth Proehl and Erin Maneri for administrative support. Finally, we thank John Gerth, Alex Keller, Andrej Krevl, Joe Little, and Jimmy Wu from Stanford IT for their close collaboration and support in conducting this study, as well as all the cybersecurity professionals who participated in our evaluation. This work is supported by the Stanford Institute for Human-Centered Artificial Intelligence Seed Grant and an unrestricted gift from OpenAI.
@@ 8 Ethics Statement
Our study was conducted with IRB approval. Participants provided informed consent to session recordings, behavioral data collection, and publication of anonymized results. They were compensated \$2,000 for their participation. Participant qualifications, including certifications and professional experience, are detailed in Appendix G. Our IRB provisions (Section 3.1) protect privacy: all data and recordings are stored on encrypted, access-controlled servers and will be deleted after the retention period. We anonymize all participants, referring to them as $P_{1}$$P_{10}$.
@@
We maintained communication with the target’s IT staff to report, triage, and patch discovered vulnerabilities per responsible disclosure. All identified vulnerabilities were remediated prior to publication. All participants read the university’s VDP and were instructed to avoid destructive actions and stay within scope. The university enforced risk-based minimum standards including monthly vulnerability management via Qualys. By collaborating with IT and recruiting security professionals as participants, we helped improve the university’s security posture. To mitigate risks from autonomous agents, a research team member monitored each session in real time with the ability to immediately terminate any out-of-scope or unsafe activity. Out-of-scope activity was defined as interaction with any systems not explicitly listed in the approved subnet scope, including other parts of the target’s infrastructure, third-party vendors and cloud providers. Unsafe activity included any actions that could cause denial-of-service conditions on production systems, permanent data corruption or deletion, account lockouts, service outages, or attempts to conduct social engineering or phishing. All monitors were computer science students with networking and security experience, and received direct training from UIT prior to the study. Termination could be enforced at three independent points: killing the agent process directly, shutting down the host machine, or severing its network connectivity. Throughout the study, we did not observe any out-of-scope or unsafe activity from any participant.